Angles of certain traingle are in AP . if the ratio of no. Of degree is least angle to the no. Of radians in the greatest base :60 /π express the angle in degree

Grade:12

1 Answers

Lab Bhattacharjee
8 years ago
\text{HINT : Let the angles be } a,a\pm d\text{ where } d\ge0 \implies a=60^\circ\\ \implies\dfrac{60}\pi=\dfrac{60-d}{(60+d)\cdot\dfrac\pi{180}}